OpenStudy (anonymous):

Isnt exactly health ... but here This isnt math... but i would really appreciate the help (: Does anyone feel up to looking over my lab and checking my journal to make sure i did it right? Tell me if you can help and i will post it all in a comment (: Would REALLY appreciate the help!!! :D

OpenStudy (anonymous):

Find an object around the house that weighs about 10 pounds, such as a heavy dictionary or a box of books. Use a bathroom scale to find the weight of your object in pounds. Next, convert the weight of the object to Newtons. One pound equals about 4.45 Newtons at Earth’s surface. Record the object you are using, its weight, and its equivalent value in Newtons in your science journal. Find a stopwatch or a kitchen timer and a meter stick. Lift the object from the floor to a height of 1 meter. Record the amount of time it takes to lift the object 1 meter. Now repeat this action 10 times in a row, as fast as you can (while being safe) and time yourself as you do so. See how many seconds it takes you to lift the object 10 times in a row, to a height of 1 meter. You might want to ask a partner to help record your time as you lift. Calculate the amount of work you did to lift the object one time. Multiply that by 10 to find out how much work you did to lift the object 10 times. Then calculate your power for lifting the object all 10 times. Report your answers in your journal. Report work in Joules and power in Watts. Please submit this journal entry as well as the journal entry from the warm-up of this section. Be sure to include both this journal entry and the one from the section warm-up. Upload and submit your completed activity for How Powerful Are

OpenStudy (anonymous):

thats my lab here is my journal so far.. please please help!?!?!?!

OpenStudy (anonymous):

44.5 newtons Object i am using: 10 lb weight Weight: 10lbs =Value in Newtons: 44.5 newtons 1. :29'11 - :29'13 2. :29'13 - :29'14 3. :29'14 - :29'15 4. :29'15 - :29'16 5. :29'16 - :29'17 6. :29'17 - :29'19 7. :29'19 - :29'20 8. :29'20 - :29'22 9. :26'22 - :26'23 10.:29'23 - :29'24 Correct Expression of Power: work / time = power Correct Expression of Work: Work = Force * Distance (Newtons = force) (distance = meters) Equation: Work = 44.5 * 1meter = 44.5 Work done when lifting once : 44.5n-m Work done when lifting 10 times: 445n-m Equation for power: power = 445 / 10 I have not finished the rest...

OpenStudy (anonymous):

lol this is funny you asked a question your self, and answered it your self. i guess it's acurate just haven't read it completely yet.